OpenStudy (anonymous):

Ok let me say this first i need to find two variables for both x and y via substitution method and only substitution method. If possible may someone answer this for me: 6x - y = 14 y = x^2 - 2

hartnn (hartnn):

what have you tried ?

OpenStudy (anonymous):

ok si will show you one sec please

OpenStudy (anonymous):

6x - x^2 - 2 = 14 -x^2 + 6x - 16 = 0 ?

OpenStudy (anonymous):

this is after substituting the equation y = x^2 -2 into 6x - y = 14

hartnn (hartnn):

y = x^2-2 -y = -x^2 +2 !

hartnn (hartnn):

so 6x - x^2 + 2 = 14 did you get how

OpenStudy (anonymous):

yeah im subbing the y in 6x - y = 14 with x^2 - 2 but after that is as far as i get

hartnn (hartnn):

i meant you wrote -2 it should be +2

OpenStudy (anonymous):

huh why?

hartnn (hartnn):

y = x^2-2 -y = ... ?

hartnn (hartnn):

6x - (x^2-2) 6x -x^2- (-2) 6x-x^2+2

OpenStudy (anonymous):

oh so 6x - x^2 + 2 = 14

hartnn (hartnn):

yep, proceed

hartnn (hartnn):

take x^2 term on right side infact every term on right side...

OpenStudy (anonymous):

ok so here is what i have so far: 6x -(x^2 - 2) = 14 6x - x^2 + 2 = 14 -x^2 + 6x - 12 = 0 (-x-3) (x -2) = 0 - this is the problem i have no clue what im supposed to do in order to get that exact variable

hartnn (hartnn):

if ab= 0 a=0 or b=0

hartnn (hartnn):

also (-x-3) (x -2) is incorrect!

OpenStudy (anonymous):

ok so what is the correct on?

OpenStudy (anonymous):

cause all i know is 2 * 3 or 4 *3 and the latter dosen't seem like the one to help me

OpenStudy (anonymous):

and that is to get -12 and the 6x with two of the same numbers

hartnn (hartnn):

are you sure the question is correct ? -x^2 + 6x - 12 = 0 does not have any real solutions!

OpenStudy (anonymous):

ok here is the variable sry i must have messed up somewhere: 6x - y = 14 y = x^2 - 2

OpenStudy (anonymous):

im ust have doen something wrong cause this is an assingment given by a teacher

hartnn (hartnn):

still same!

hartnn (hartnn):

or you can show that -x^2 + 6x - 12 = 0 does not have any real solutions and hence no real solution exist for the system

OpenStudy (anonymous):

oh ok thank you well that is all for today thank you very much

hartnn (hartnn):

welcome ^_^